# Robust Energy-Efficient DRL-Based Optimization in UAV-Mounted RIS Systems with Jitter

### v1

@@ II System Model
We consider the UAV-mounted RIS wireless communication system illustrated in Fig. 1, where a ground base station (BS) located at $\mathbf{q}_{\mathrm{BS}}=[0,0,H_{\mathrm{BS}}]^{T}$ is equipped with $M=M_{y}M_{z}$ antennas, spaced by $\Delta_{y}^{\mathrm{BS}}$ and $\Delta_{z}^{\mathrm{BS}}$ along the horizontal and vertical axes, respectively. The BS serves $K$ single-antenna ground users with locations given by $\mathbf{q}_{\mathrm{U}_{k}}=[x_{k},y_{k},0]^{T}$, $k=1,\dots,K$ assisted by a UAV-mounted RIS, which is implemented as a uniform planar array with $N=N_{x}N_{y}$ passive reflecting elements. The UAV hovers at a fixed altitude $H_{\mathrm{RIS}}$, with center coordinate $\mathbf{q}_{\mathrm{RIS}}=[x_{\mathrm{RIS}},y_{\mathrm{RIS}},H_{\mathrm{RIS}}]^{T}$. Each RIS element is uniformly spaced with inter-element distances $\Delta_{x}^{\mathrm{RIS}}$ and $\Delta_{y}^{\mathrm{RIS}}$ along the planar array. We assume far-field propagation for both BS–RIS and RIS–$U_{k}$ links, enabling angular modeling via azimuth and elevation angles [11]. The angle of departure (AoD) from the BS is denoted as $(\phi_{t,\mathcal{G}},\vartheta_{t,\mathcal{G}})$, and the angle of arrival (AoA) at the RIS is $(\phi_{r,\mathcal{G}},\vartheta_{r,\mathcal{G}})$. The AoD from the RIS to the $k$-th user is given by $(\phi_{t,\mathcal{H}},\vartheta_{t,\mathcal{H}})$. The transmit array response at the BS is modeled as the Kronecker product of its vertical and horizontal steering vectors with respect to the AoD as
@@ II-A Channel Model
We adopt an air-to-ground (ATG) channel model for the BS–RIS link [14, 11], and a Rician fading model for the RIS–$U_{k}$ link [8]. The BS-to-RIS channel matrix $\mathbf{G}\in\mathbb{C}^{N\times M}$ comprises both LoS and non-line-of-sight (NLoS) components, modeled as
@@ II-D UAV-mounted RIS Jitter Modeling
[…]
@@ III Problem Formulation
s.t. $R_{k}(t)\geq R_{\min},\quad\forall k\in\mathbb{K},\ t\in\mathbb{T},$ (21b)
$0\leq\tau(t)\leq 1,\quad\forall t\in\mathbb{T},$ (21c)
$0\leq P_{BS}(t)\leq P_{BS}^{\max},\quad\forall t\in\mathbb{T},$ (21d)
$0\leq p_{k}(t)\leq P_{U}^{\max},\quad\forall k\in\mathbb{K},\ t\in\mathbb{T},$ (21e)
$\theta_{n}(t)\in[0,2\pi],\quad\forall n\in[1,N],\ t\in\mathbb{T},$ (21f)
$|e^{j\theta_{n}(t)}|=1,\quad\forall n\in[1,N],\ t\in\mathbb{T},$ (21g)
@@ V Simulation Results
$A_{\text{sim}}$ Simulation area $20$ m $\times$ $20$ m
$N$ Number of RIS elements $16$
@@
$\beta_{0}$ Reference path loss at $d_{\text{ref}}=1$ m $-30$ dB
$\sigma_{k}^{2}$ Noise power at user $k$ $-102$ dBm

### v2

@@ II System Model
[…]
@@ II-A Channel Model
[…]
@@ II-D UAV-mounted RIS Jitter Modeling
where $\mathbf{e}_{x}$ and $\mathbf{e}_{z}$ are unit vectors along the $x$- and $z$-axes, $\mathcal{P}_{xy}(\cdot)$ denotes the orthogonal projection onto the xyxy-plane, and $\langle\cdot,\cdot\rangle$ is the Euclidean inner product. These perturbed angles are then used to update the channel components in (3) and (5), thereby incorporating the impact of UAV jitter.
@@ III Problem Formulation
s.t. $R_{k}(t)\geq R_{\min},\qquad\forall k\in\mathbb{K},\ t\in\mathbb{T},$ (21b)
$0\leq\tau(t)\leq 1,\qquad\forall t\in\mathbb{T},$ (21c)
$0\leq P_{BS}(t)\leq P_{BS}^{\max},\qquad\forall t\in\mathbb{T},$ (21d)
$0\leq p_{k}(t)\leq P_{U}^{\max},\qquad\forall k\in\mathbb{K},\ t\in\mathbb{T},$ (21e)
$\theta_{n}(t)\in[0,2\pi],\qquad\forall n\in[1,N],\ t\in\mathbb{T},$ (21f)
$|e^{j\theta_{n}(t)}|=1,\qquad\forall n\in[1,N],\ t\in\mathbb{T},$ (21g)
@@ V Simulation Results
$A_{\text{sim}}$ Simulation area 2020 m $\times$ 2020 m
NN Number of RIS elements $16$
@@
$\beta_{0}$ Reference path loss at $d_{\text{ref}}=1$ m 30-30 dB
$\sigma_{k}^{2}$ Noise power at user $k$ 102-102 dBm